OpenStudy (anonymous):

find each specific vector or scalar. (1) v+w (2) u-v

OpenStudy (danjs):

if those are vectors, addition of 2 vectors results in another vector

OpenStudy (danjs):

for the second, you add the negative of v, v is flipped aound 180 degrees first

OpenStudy (anonymous):

I really don't understand this at all because it has no numbers how would i find the answer

OpenStudy (danjs):

something is missing in the prob you gave i think, is this just assuming there are 3 vectors, u,v,and w?

OpenStudy (anonymous):

without knowing \(u\) or \(v\) it is not possible

OpenStudy (danjs):

if v + w = u then you can figure out them, is there a diagram

OpenStudy (anonymous):

i am going to try and seen shot it from my online book

OpenStudy (anonymous):

it is number 22 and 23

OpenStudy (danjs):

lol, look at the directions for that section of probs

OpenStudy (danjs):

They gave you all three of the vectors components

OpenStudy (anonymous):

okay yes i see that now but how do i apply that?

OpenStudy (danjs):

Each vector has components \[\large u = <u _{x}, u _{y}> ~~~~or~~~~u = u _{x}*i + u _{y}*j\]

OpenStudy (danjs):

like a triangle in 2-space plane, a line will have a component parallel to each axis

OpenStudy (danjs):

8right triangle

OpenStudy (danjs):

For vector addition, you can simply add the like components to get a new vector look at the definition

OpenStudy (anonymous):

for the first one it would be -4+j?

OpenStudy (anonymous):

-4i*

OpenStudy (danjs):



OpenStudy (anonymous):

okay when you said combine like components it would be -4i+j right?

OpenStudy (danjs):

vector addition u = a*i + b*j v = c*i + d*j u + v = (a + c)i + (b + d)j

OpenStudy (danjs):

you sum each component for i,j,k

OpenStudy (anonymous):

so i did it right?

OpenStudy (danjs):

v + w = < -3 + (-1) , 7 + (-6) >

OpenStudy (danjs):

yep, -4i +1j, or <-4,1>

OpenStudy (anonymous):

okay great now what do i use that for?

OpenStudy (anonymous):

is that the vector?

OpenStudy (anonymous):

thank you so much

OpenStudy (danjs):

that is it, just a vector sum

OpenStudy (danjs):

yes, that results in a new vector